OpenStudy (anonymous):

Write the equation for the following situation: In one U.S. town, the annual consumption, b, of beef(in kg per person) can be estimated by adding 20 to half the years, t, since 1975.

OpenStudy (anonymous):

If t is the given year, then number of years since 1975 is (t - 1975). Half of that is (t - 1975)/2 Thus, the equation is : b = 20 + (t - 1975)/2 If t = 1989, then b = 20 + (1989 - 1975)/2 = 20 + 14/2 = 27 kg.

OpenStudy (anonymous):

If t is the given year, then number of years since 1975 is (t - 1975). Half of that is (t - 1975)/2 Thus, the equation is : b = 20 + (t - 1975)/2 If t = 1989, then b = 20 + (1989 - 1975)/2 = 20 + 14/2 = 27 kg. kumar 2006 is correct